OpenStudy (anonymous):

The demand and cost functions for a product are p(x) = 110 − 0.01x C(x) = 30x + 1400 x = # of units produced weekly If the manufacturer decides to increase production by 70 units per week, find the rate at which profit is changing with respect to time when the weekly production is 3600 units.

OpenStudy (perl):

thats a profit equation p(x) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

That's the demand function

OpenStudy (perl):

then p stands for price

OpenStudy (perl):

p is price per unit multiply that by x , number of units produced/sold , to get revenue. then profit = revenue - cost

OpenStudy (anonymous):

So revenue is R(x) = 110x-0.01x^2

OpenStudy (anonymous):

and profit is \[P(x) = 80x - 0.01x^2-1400\]

OpenStudy (anonymous):

Now do I differentiate P(x)?

OpenStudy (perl):

yes, but i got a different function

OpenStudy (perl):

one moment

OpenStudy (perl):

sorry you are right

OpenStudy (anonymous):

Ok so I got

OpenStudy (anonymous):

\[\frac{ dP }{ dt } = 80 - 0.01x(\frac{ dx }{ dt }\]

OpenStudy (anonymous):

Oh i did that wrong

OpenStudy (anonymous):

So P'(x) = 80-0.02x

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

\[P'(x) = 80-0.02x(\frac{ dx }{ dt })\]

OpenStudy (perl):

$$ \Large { \Large{ Profit =Revenue - Cost \\ \\ \therefore \\ P(x) = R(x) - C(x) \\ P(x) = x* p(x) - C(x) \\P(x) = x (110 - 0.01x) - (30x + 1400) \\P(x) = 80*x-0.01*x^2-1400 \\ \therefore\\ \frac{dP}{dt} = 80\cdot \frac{dx}{dt} - (0.01)\cdot 2x \cdot \frac{dx}{dt} } } $$

OpenStudy (anonymous):

After this, what do I do? Plug in 70 for dx/dt and 3600 for x?

OpenStudy (anonymous):

Ok I did that and got the answer

OpenStudy (perl):

is it correct?

OpenStudy (anonymous):

Yes :^)

OpenStudy (perl):

$$ \Large { \Large{ Profit =Revenue - Cost \\ \\ \therefore \\ P(x) = R(x) - C(x) \\ P(x) = x* p(x) - C(x) \\P(x) = x (110 - 0.01x) - (30x + 1400) \\P(x) = 80*x-0.01*x^2-1400 \\ \therefore\\ \frac{dP}{dt} = \frac{dP}{dx} \cdot \frac{dx}{dt} = (80 - 0.01\cdot 2x) \cdot \frac{dx}{dt} } } $$